OpenStudy (anonymous):

if 1.7 plus 4.5 times a number equals the sum of a number and 61.4, whats the number

OpenStudy (anonymous):

17.057

OpenStudy (anonymous):

\[1.7+4.5x=x+61.4\]\[3.5x=61.4-1.7\]\[x=\frac{61.4-1.7}{3.5}\]\[x=17.05714\]

OpenStudy (anonymous):

can 17.057 rounded up

OpenStudy (anonymous):

17.1